HUBBLE SPACE TELESCOPE DAILY REPORT #1722 PERIOD COVERED: 0000Z (UTC) 09/19/96 - 0000Z (UTC) 09/20/96 Daily Status Report as of 264/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: = Decontamination) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 1.2 Completed FOS/BL 6831 (High Speed Photometry of the UV Counterpart = to 4U1820-30) The Faint Object Spectrograph (Blue Side) was used to make Rapid = Mode, ultraviolet, spectrographic observations of the counterpart of = x-ray, binary object 4U1820-30. The target is a low mass binary system = with an orbital period of 11 minutes. Approximately fifteen minutes = after this 8-hour proposal began, at 263/015112Z, there were two FOS = status buffer messages that indicated that the mode 2 target acquisition = had failed. HSTAR 5716 was written. Otherwise, the observations were = executed as scheduled, and no further anomalies were noted. 1.3 Completed WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 1.4 Completed FOS/RD 6696 (Studying the UV Emission line Spectrum of = High Z radio Galaxies) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet spectrographic observations of the extended line regions in = the high Z, radio galaxy PKS0349-27. Several different locations within = the object were observed. The observations were completed as planned, = and no problems were reported. 1.5 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.6 Completed WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with = the prime science observations as part of the Medium Deep Sky Survey. = The observations were executed as scheduled, and no problems were = reported. 1.7 Completed GHRS/2 6078 (The Wind Momentum-Luminosity Relationship = for SMC A and B Supergiants.) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the O/B class star AV683 which IS located = in the Small Magellanic Cloud in order to measure the wind = momentum-luminosity relationship in this low metallicity stars. This = relationship is used to determine the distance to galaxies. The = observations were completed as planned and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 The guide star acquisition at 264/071300Z defaulted to fine lock = backup on one FGS only. HSTAR 5717 was written. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: There were no reported operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH